# Trustworthy Retrosynthesis: Eliminating Hallucinations with a Diverse Ensemble of Reaction Scorers

### v2

@@ 1 Introduction
The advent of deep generative modeling has transformed a broad range of domains, and resulted in impressive applications such as photorealistic image synthesis, code generation, and automated theorem proving (Rombach et al., 2022; Li et al., 2022; Yang et al., 2023). The undisputed efficacy of generative models is nonetheless hindered by the possibility of factually wrong or outright nonsensical output, often called "hallucinations" Sahoo et al. (2024). In some applications, such as automated theorem proving, hallucinations can be curtailed with formal verification of output. In others, they can only be mitigated through partial and often imprecise verification.
Retrosynthesis, the task of constructing synthetic routes — sequences of chemical reactions that lead to a desired target molecule from simpler precursors — is another domain that has undergone significant developments with the rise of generative modeling (Coley et al., 2018). Modern systems typically generate complete synthetic routes by iteratively applying predictions from a single-step retrosynthesis (SSR) model within a graph-based search algorithm. When reactions have no known chemical precedent or display features such as unstable reagents, they jeopardize the validity of the synthetic path they are part of. Generative SSR models inherit the limitations of generative modeling: the reactions they propose are frequently hallucinated, that is, chemically unsound (Torren-Peraire et al., 2024). An example of a hallucinated reaction is shown in Figure 1
@@
In this work, we propose RetroTrim, a retrosynthesis system designed to tackle the issue of hallucinated reactions by pairing a performant SSR generator with an ensemble of complementary reaction scorers as an in-scope filter. We show that RetroTrim is the only among common retrosynthetis solutions that successfully avoids all hallucinated reactions in proposed synthetic plans for a set of unpublished challenging drug-like targets. Moreover, RetroTrim does so while leading in the number of targets for which a synthetic plan without issues is found. This approach proved highly effective in practice, securing first place in the Standard Industries Retrosynthesis Challenge, in which the correctness of pathways judged by human experts was crucial.
Our approach in developing RetroTrim is based on the principle of ensemble learning Hansen & Salamon (1990): combining diverse scorers with distinct error patterns yields a more robust and accurate assessment of reaction plausibility. The ensemble scorer (meta-scorer) functions as an external validator of the generative SSR model. It prunes all reactions below a given score threshold from the synthesis tree, curtailing expansion of hallucinated nodes. For an overview see Fig. 2. The three scorers which make up the in-scope filter are as follows:
@@
    Reaction Prior (RP): A Transformer-based architecture Vaswani et al. (2017) whose scoring function is designed to mimic the considerations chemists take into account when evaluating reaction plausibility. Its training and scoring method lends itself to discovering a broad spectrum of hallucinations.
@@
We compare performances of RP, RGP and RRS on a dataset of retrosynthetic intermediates. We show that each scorer excels at filtering different kinds of hallucinations, and that the meta-scorer further improves performance in most areas.
To measure the performance of both retrosynthesis systems as well as reaction scorers, we employ a novel reaction validation protocol in which PhD-level chemists assess each reaction in a given synthetic plan according to a predefined labeling schema. Reactions are sorted into seven categories depending on the kind of issue they raise: Magic, Selectivity, Functional group incompatibility, Reactivity, One pot, Unstable, or Reactants mismatch; and according to the severity of said issue: Worthwhile, Rather not, and Nonsense. Reactions which present no issues are given the special labels No problem and Safe bet, respectively. Under this categorization, Nonsense are the reactions that correspond to hallucinations, while Worthwhile and Rather not reactions contain mild or moderate issues which makes them undesirable or unpredictable. The overall confidence in a synthetic pathway is determined by the reaction with the highest issue severity within it.
All retrosynthesis systems are evaluated on a set of thirty-two challenging drug-like targets. These targets were defined so as to match common structural features present in modern drugs while not being close analogs of any structure found in synthetic literature. This ensures no data leakage occurred in-between training and inference of compared models, while still being representative of the medicinal chemistry domain.
@@
Figure 2: Visualization of RetroTrim. The generator proposes precursor molecules for a given target, which are then evaluated for reaction plausibility. Plausible reactions proceed to the next retrosynthetic step for non-commercially available precursors (blue boxes), while implausible reactions terminate the branch. The search concludes when a complete synthetic route from market-available starting materials (purple boxes) to the target molecule is found. The displayed reactions originate from retrosynthetic pathways generated by our systems and the baselines. The "Safe Bet" route is strongly supported by literature with multiple documented precedents, while the "Nonsense" route lacks any published examples.
@@ 2.1 Multi-Step Retrosynthesis
Multi-step retrosynthesis constructs complete synthetic routes by iteratively applying single-step retrosynthesis (SSR) predictions. Classical search strategies, such as Monte Carlo tree search Segler et al. (2018) and A* search (Retro*) Chen et al. (2020), use SSR models to expand compounds during generation of synthetic pathways.
SSR models can be categorized into template-based, template-free, and semi-template-based approaches. Template-based methods (e.g., RetroSim (Coley et al., 2017), NeuralSym (Segler & Waller, 2017), GLN (Dai et al., 2019a)) use predefined reaction templates, ensuring interpretability, but their coverage is limited and their construction involves trade-offs that can lead to invalid template application. Template-free models (e.g., sequence-to-sequence Transformers (Zhong et al., 2022), GNNs (Liu et al., 2017; Karpov et al., 2019; Sacha et al., 2021)) learn chemical rules directly from data, offering scalability and the ability to generalize to novel reactions, but with little guarantee as to the validity of output. Semi-template-based approaches (e.g., GraphRetro (Somnath et al., 2021), RetroXpert (Yan et al., 2020)) combine templates with learned representations, inheriting both the advantages and limitations of both approaches. Regardless of their category, SSR models are all prone to erroneous output Coley et al. (2018).
@@
To address this, round-trip accuracy (Schwaller et al., 2020) is commonly used. This metric evaluates a predicted reaction by passing the predicted reactants through a forward reaction model and checking if the original product is recovered. Round-trip accuracy is considered a better proxy for chemical validity than top-k accuracy. However, it remains unclear how well round-trip accuracy reflects the actual plausibility of the predicted reactions. Currently, there are no exhaustive data quantifying how many predictions that pass or fail the round-trip check correspond to chemically plausible reactions.
Evaluating multi-step retrosynthetic routes is typically limited to basic pathway properties, such as the number of steps, route length, branching factor, or overall synthesizability (Maziarz et al., 2025). These metrics provide little insight into the actual correctness or chemical plausibility of the complete synthetic route.
@@ 2.2 Mid-Search Reaction Validation
The plausibility of multi-step retrosynthetic routes has drawn increasing attention from researchers, leading to the development of methods aimed at improving pathway validity. A commonly used approach is to employ a feasibility model as a filter. Such models are typically trained as classifiers to distinguish correct reactions from artificially generated negative examples, which are often created by perturbing positive reactions, for example by randomly modifying substrates or swapping the original product with a chemically similar one (Segler et al., 2018; Genheden et al., 2020a). A forward model (predicting products based on substrates) trained only on positive data was used in the same way in IBM RXN Schwaller et al. (2020).
Another strategy is to use the likelihood of a generated reaction as a scoring function during search. This can be computed from template-free models, such as the Molecular Transformer (Schwaller et al., 2019) architecture, where the model’s confidence, derived from output token probabilities, has been shown to correlate with reaction correctness, or from template-based models using softmax scores, as in RetroFallback (Tripp et al., 2023). In order to improve the quality of predicted reactions, RetroGFN (Gaiński et al., 2024) was trained using feedback from a forward model and round-trip accuracy. RetroChimera (Maziarz et al., 2024) takes a complementary approach by enforcing reaction validity through selecting the most plausible predictions from an ensemble of diverse reaction generators, effectively combining their strengths to increase overall accuracy.
Finally, evidence-based validation via retrieval grounds model predictions in established chemical knowledge, mirroring a chemist’s workflow of searching for literature precedents. Retrieval-augmented methods, such as the Retrieval-Augmented RetroBridge (RARB) framework, retrieve similar molecules from a database to guide the generation of reactants (Qiao et al., 2025). These approaches help ensure that generated reactions are consistent with known chemistry, further improving the plausibility of multi-step routes. In the context of these works, RetroTrim is the first to use retrieved reactions to filter predictions of a reaction generation.
@@ 3 Methods
RetroTrim is built around a central, best-performing scorer called Reaction Prior, supported by two additional scorers designed to compliment it by targeting specific error types. These three scorers are aggregated into a Meta-Scorer that provides the final, robust assessment of reaction correctness.
@@ 3.1 Reaction Prior
The Reaction Prior (RP) is a novel method inspired by how experienced chemist reason about reactions: considering the reaction globally, assessing the reaction center, and comparing it to alternative reactions that could occur. The RP score $S_{final}$ is thus a weighted combination of three components: $S_{final}=S_{RP}^{\alpha}\cdot S_{Regio}^{\beta}\cdot S_{RC}^{\gamma}$. Here, $S_{RP}$ is the Reaction Prior (global) Score, $S_{RC}$ is the Reaction Center Score, and $S_{Regio}$ is the Regioselectivity Score, with $\alpha$, $\beta$, and $\gamma$ serving as weighting factors. RP is implemented as an autoregressive, encoder-decoder BART (Lewis et al., 2019) architecture, where substrates and products are processed by the decoder, trained for next-token prediction by minimizing cross-entropy loss.
@@
This score reflects the overall feasibility of the reaction. It is the log probability assigned by the model to the reaction sequence, normalized by the square root of the total number of tokens ($T$): $S_{RP}=\frac{1}{\sqrt{T}}\log P(\text{reaction})$.
@@
This score evaluates the model’s confidence in the identified reactive sites. It is the sum of log probabilities for the tokens representing atoms at the reaction’s core, normalized by the number of such tokens ($T_{RC}$): $S_{RC}=\frac{1}{T_{RC}}\sum_{i\in\text{reaction center}}\log P(\text{token}_{i})$.
@@
This component quantifies reaction site specificity. It is calculated by comparing the probability of the desired reaction ($P_{\text{desired}}$) to the summed probabilities of all alternative reactions at different potential sites ($P_{\text{undesired}}$): $S_{Regio}=\log\left(\frac{P_{\text{desired}}}{P_{\text{undesired}}+\epsilon}\right)$, where $\epsilon$ is a small constant to prevent division by zero.
@@ 3.2 Reaction Graph Plausibility
A Graph Attention Network (GAT) (Veličković et al., 2017) is trained to differentiate chemically valid reactions from implausible ones. Training uses reaction datasets for positive examples and synthetic negative examples generated through forward and two-step backward template applications. This approach is similar to those proposed by (Segler et al., 2018; Genheden et al., 2020a), but it uses a graph neural network instead of a feedforward network with fingerprint inputs, and employs more sophisticated artificial negative reactions, generated not only by applying random templates in the forward direction but also through retro-synthetic random template applications, which increases the diversity of types of generated incorrect reactions. Details of GAT featurization are described in Appendix B.
@@ 3.3 Reference Reaction Scorer
Our Reference Reaction Retrieval Scorer (RRS) quantifies reaction plausibility through a logarithmic transformation of the unique reference reaction count within the candidate reaction’s coarse-grained cluster and fine-grained cluster:
$p(reaction)=\log(n_{ref}(reaction)+1)$ (1)
where $n_{ref}(reaction)$ represents the unique number of reference reactions in the coarse-grained and fine-grained clusters containing reactionreaction.
3.4 Meta-Scorer Aggregation
To improve reaction filtering, our Meta-Scorer integrates scores from Reaction Prior (RP) and RGP models, and empirical precedents ($n_{\text{ref}}>0$) retrieved via the pipeline in Sec. 3.3. This hybrid approach mitigates the weaknesses of purely data-driven or precedent-based methods. The continuous score is described by equation $\text{score}_{\text{META}}=\max(\text{score}_{\text{RGP}},\text{score}_{\text{RP}})$ if $n_{\text{ref}}>0$ (0 otherwise).
For binary classification tasks and search, reactions are filtered using predefined thresholds, selected through grid search to balance the recall and precision:
$\text{score}_{\text{META}}=\begin{cases}1&\text{if }\text{score}_{\text{RGP}}>thr_{RGP}\text{ and }\text{score}_{\text{RP}}>thr_{RP}\text{ and }n_{\text{ref}}>0\\ 0&\text{otherwise}\end{cases}$ (2)
By synthesizing diverse evidence types Meta-Scorer enables more reliable reaction filtering for multi-step synthesis planning. This integrated approach mitigates individual weaknesses of purely data-driven or precedent-based methods, yielding improved performance.
3.5 Integration with Search (Retro*)
The calibrated Meta-Scorer is used during multi-step retrosynthesis search to improve the quality of the pathways predicted by the BART generator. We integrate the scorer into the Retro* (Chen et al., 2020) search framework as a reaction filtering mechanism. Reactions are pruned from the search tree if $\text{score}_{\text{META}}$ defined in 2 is equal to 0.

### v3

@@ 1 Introduction
The advent of deep generative modeling has transformed a broad range of domains, and resulted in impressive applications such as photorealistic image synthesis, code generation, and automated theorem proving (RealisticImageGen; AlphaCode; LeanDojo). The undisputed efficacy of generative models is nonetheless hindered by the possibility of factually wrong or outright nonsensical output, often called "hallucinations" (HallucinationsSurvey). In some applications, such as automated theorem proving, hallucinations can be curtailed with formal verification of output. In others, they can only be mitigated through partial and often imprecise verification.
Retrosynthesis, the task of constructing synthetic routes — sequences of chemical reactions that lead to a desired target molecule from simpler precursors — is another domain that has undergone significant developments with the rise of generative modeling (Coley2018MachineLI)). Modern systems typically generate complete synthetic routes by iteratively applying predictions from a single-step retrosynthesis (SSR) model within a graph-based search algorithm. When reactions have no known chemical precedent or display features such as unstable reagents, they jeopardize the validity of the synthetic path they are part of. Generative SSR models inherit the limitations of generative modeling: the reactions they propose are frequently hallucinated, that is, chemically unsound (Peraire2024). An example of a hallucinated reaction is shown in Figure 1
@@
In this work, we propose RetroTrim, a retrosynthesis system designed to tackle the issue of hallucinated reactions by pairing a performant SSR generator with an ensemble of complementary reaction scorers as a plausibility filter. We show that RetroTrim is the only among common retrosynthetis solutions that successfully avoids all hallucinated reactions in proposed synthetic plans for a set of unpublished challenging drug-like targets. Moreover, RetroTrim does so while leading in the number of targets for which a synthetic plan without issues is found. This approach proved highly effective in practice, securing first place in the Standard Industries Retrosynthesis Challenge, in which the correctness of pathways judged by human experts was crucial.
Our approach in developing RetroTrim is based on the principle of ensemble learning (ensembles): combining diverse scorers with distinct error patterns yields a more robust and accurate assessment of reaction plausibility. The ensemble scorer (MetaScorer) functions as an external validator of the generative SSR model. It prunes all reactions below a given score threshold from the synthesis tree, curtailing expansion of hallucinated nodes. For an overview see Fig. 2. The three scorers which make up the plausibility filter are as follows:
@@
    Reaction Prior (RP): A Transformer-based architecture (aTTeNtiONisAllYOunEEd) whose scoring function is designed to mimic the considerations chemists take into account when evaluating reaction plausibility. Its training and scoring method lends itself to discovering a broad spectrum of hallucinations.
@@
We compare performances of RP, RGP and RRS on a dataset of retrosynthetic intermediates. We show that each scorer excels at filtering different kinds of hallucinations, and that the MetaScorer further improves performance in most areas.
In identifying hallucinations of retrosynthesis systems, common automated metrics can at best serve as surrogates: algorithmic procedures cannot model the whole concept of chemical plausibility, and exact matches to literature are rare in practice for complex targets. The only way to arrive at a verdict on a synthetic route in practice is to propose it to one or more expert chemists. If the route contains reactions for which no known synthetic methodology exists, such a route will be rejected without the need for experimental verification.
For this reason, to measure the performance of both retrosynthesis systems as well as reaction scorers, we employ a novel reaction validation protocol in which PhD-level chemists assess each reaction in a given synthetic plan according to a predefined labeling schema. Reactions are sorted into seven categories depending on the kind of issue they raise: Magic, Selectivity, Functional group incompatibility, Reactivity, One pot, Unstable, or Reactants mismatch; and according to the severity of said issue: Worthwhile, Rather not, and Nonsense. Reactions which present no issues are given the special labels No problem and Safe bet, respectively. Under this categorization, Nonsense corresponds to hallucinated reactions, while Worthwhile and Rather not reactions contain mild or moderate issues which make them undesirable or unpredictable. The overall confidence in a synthetic pathway is determined by the reaction with the highest issue severity within it.
All retrosynthesis systems are evaluated on a set of thirty-two challenging drug-like targets (see Appendix D). These targets were defined so as to match common structural features present in modern drugs while not being close analogs of any structure found in synthetic literature. This ensures no data leakage occurred in-between training and inference of compared models, while still being representative of the medicinal chemistry domain.
@@
Figure 2: Visualization of RetroTrim (above) and retrosynthetic search with plausibility filtering (below). RetroTrim encompasses a generator, which proposes precursor molecules for a given target, and a scorer, which evaluates the plausibility of the generated reaction. In the search process, plausible precursors are expanded further, until we arrive at commercially available starting materials. Implausible reactions terminate the search branch. The search concludes when a complete synthetic route from commercially-available starting materials to the target molecule is found.
@@ 2.1 Multi-Step Retrosynthesis
Multi-step retrosynthesis constructs complete synthetic routes by iteratively applying single-step retrosynthesis (SSR) predictions. Classical search strategies, such as Monte Carlo tree search (Segler2018) and A* search (Retro*) (RetroStar), use SSR models to expand compounds during generation of synthetic pathways.
SSR models can be categorized into template-based, template-free, and semi-template-based approaches. Template-based methods (e.g., RetroSim (ColeyRetroSim), NeuralSym (SeglerNeuralSym), GLN (DaiGLN)) use predefined reaction templates, ensuring interpretability, but their coverage is limited and their construction involves trade-offs that can lead to invalid template application. Template-free models (e.g., sequence-to-sequence Transformers (zhong2022rootaligned), GNNs (LiuExpertSys; KarpovTransformer; SachaMEGAN)) learn chemical rules directly from data, offering scalability and the ability to generalize to novel reactions, but with little guarantee as to the validity of output. Semi-template-based approaches (e.g., GraphRetro (SomnathGraphRetro), RetroXpert (YanRetroeXpert)) combine templates with learned representations, inheriting both the advantages and limitations of both approaches. Regardless of their category, SSR models are all prone to erroneous output (Coley2018MachineLI).
@@
To address this, round-trip accuracy (SchwallerIBMRXN) is commonly used. This metric evaluates a predicted reaction by passing the predicted reactants through a forward reaction model and checking if the original product is recovered. Round-trip accuracy is considered a better proxy for chemical validity than top-k accuracy.
However, it remains unclear how well round-trip accuracy reflects the actual plausibility of the predicted reactions. There is currently no exhaustive data quantifying how many predictions that pass or fail the round-trip check correspond to chemically plausible reactions.
Evaluating multi-step retrosynthetic routes is typically limited to basic pathway properties, such as the number of steps, route length, branching factor, or overall synthesizability (MaziarzSyntheseus). These metrics provide little insight into the actual correctness or chemical plausibility of the complete synthetic route.
Evaluation of multi-step routes by expert chemists provides a more trustworthy viability assessment. To our knowledge, the only work containing human evaluation of routes powered by generative models is Segler2018, which conducts a double-blind A/B test comparing routes existing in literature and those produced by their model. Their results showed that for the first time, computer-generated routes could be on-par or preferred to those found in the literature by chemists. While a significant result in its own right, their test only recorded preference between paths as opposed to measuring the rate of significant system errors. It also spanned only 9 randomly chosen targets. In contrast, we conduct fine-grained error-analysis of routes proposed by 4 modern retrosynthesis systems, as well as 6 variants of our own system, including the full RetroTrim and simpler baselines. Moreover, we do so on a set of 32 targets chosen for their challenging nature.
@@ 2.2 Mid-Search Reaction Validation
The plausibility of multi-step retrosynthetic routes has drawn increasing attention from researchers, leading to the development of methods aimed at improving pathway validity. A commonly used approach is to employ a plausibility model as a filter. Such models are typically trained as classifiers to distinguish correct reactions from artificially generated negative examples, which are often created by perturbing positive reactions, for example by randomly modifying substrates or swapping the original product with a chemically similar one (Segler2018; aizynthfinderScorer). A forward model (predicting products based on substrates) trained only on positive data was used in the same way in IBM RXN (SchwallerIBMRXN).
Another strategy is to use the likelihood of a generated reaction as a scoring function during search. This can be computed from template-free models, such as the Molecular Transformer (SchwallerTransformerUncertainty) architecture, where the model’s confidence, derived from output token probabilities, has been shown to correlate with reaction correctness, or from template-based models using softmax scores, as in RetroFallback (retrofallback). In order to improve the quality of predicted reactions, RetroGFN (gainski2024retrogfn) was trained using feedback from a forward model and round-trip accuracy. RetroChimera (maziarz2024chimera) takes a complementary approach by selecting the most plausible predictions from an ensemble of diverse reaction generators, effectively combining their strengths to increase overall accuracy.
Finally, evidence-based validation via retrieval grounds model predictions in established chemical knowledge, mirroring a chemist’s workflow of searching for literature precedents. Retrieval-augmented methods, such as the Retrieval-Augmented RetroBridge (RARB) framework, retrieve similar molecules from a database to guide the generation of reactants (QiaoRARB). These approaches help ensure that generated reactions are consistent with known chemistry, further improving the plausibility of multi-step routes. In the context of these works, RetroTrim is the first to use retrieved reactions to filter predictions of a reaction generation.
@@ 3 Methods
RetroTrim is a combination of an SSR reaction generator and a reaction scorer. The key innovation in RetroTrim is the design of the scorer, which is built around a central, best-performing scorer called Reaction Prior, supported by two additional scorers: Reaction Graph Plausibility scorer and Reference Reaction Scorer. The additional scorers are designed to compliment Reaction Prior by targeting specific error types. Together, they aggregate into the MetaScorer which provides the final, robust assessment of reaction correctness.
@@ 3.1 Reaction Prior
The Reaction Prior (RP) is a novel method inspired by how experienced chemist reason about reactions: considering the reaction globally, assessing the reaction center, and comparing it to alternative reactions that could occur. The RP score $S_{final}$ is thus a weighted combination of three components: $S_{final}=S_{RP}^{\alpha}\cdot S_{Regio}^{\beta}\cdot S_{RC}^{\gamma}$. Here, $S_{RP}$ is the Reaction Prior (global) Score, $S_{RC}$ is the Reaction Center Score, and $S_{Regio}$ is the Regioselectivity Score, with $\alpha$, $\beta$, and $\gamma$ serving as weighting factors. These weights can be tuned to optimize a metric of interest. In practice, we find $\alpha=1$, $\beta=1.5$ and $\gamma=2.5$ to be reasonable default values for drug-like targets. RP is implemented as an autoregressive, encoder-decoder BART (LewisBART) architecture, where substrates and products are processed by the decoder, trained for next-token prediction by minimizing cross-entropy loss.
@@
This score reflects the overall feasibility of the reaction. It is proportional to the log probability assigned by the model to the reaction sequence. Analogously to language models, which assign higher probabilities to sequences close to the training set, this score acts as a measure of reaction similarity to the dataset on which the model is trained on. In this respect, it mimics the way in which a chemist would look for precedent for the overall transformation. We normalize the log probability by the square root of the total number of tokens ($T$), which we find to maximize predictive performance in practice: $S_{RP}=\frac{1}{\sqrt{T}}\log P(\text{reaction})$.
@@
When evaluating reactions, chemists pay special attention to the sites where changes in connectivity occur. The $S_{RC}$ score analogously evaluates the model’s confidence in the identified reactive sites. It is proportional to the sum of log probabilities of tokens representing atoms in the reaction’s center. Unlike $S_{RP}$, here we normalize by the number of such tokens ($T_{RC}$), similarly guided by empirical calibration: $S_{RC}=\frac{1}{T_{RC}}\sum_{i\in\text{reaction center}}\log P(\text{token}_{i})$.
@@
This component quantifies reaction site specificity. It is calculated by comparing the probability of the reaction at the given reaction site ($P_{\text{desired}}$) to the summed probabilities of all sites where the reaction might occur ($P_{\text{undesired}}$): $S_{Regio}=\log\left(\frac{P_{\text{desired}}}{P_{\text{undesired}}+\epsilon}\right)$, where $\epsilon$ is a small constant to prevent division by zero. This score component reflects the tendency for chemists to evaluate whether the particular site where a reaction occurs is preferred compared to potential alternatives.
@@ 3.2 Reaction Graph Plausibility
A Graph Attention Network (GAT) (GAT) is trained to differentiate chemically valid reactions from implausible ones. Training uses reaction datasets for positive examples and synthetic negative examples generated through forward and two-step backward template applications. This approach is similar to those proposed by (Segler2018; aizynthfinderScorer), but it uses a graph neural network instead of a feedforward network with fingerprint inputs, and employs more sophisticated artificial negative reactions, generated not only by applying random templates in the forward direction but also through retro-synthetic random template applications, which increases the diversity of types of generated incorrect reactions. Details of GAT featurization are described in Appendix B.
@@ 3.3 Reference Reaction Scorer
Our Reference Reaction Retrieval Scorer (RRS) quantifies reaction plausibility through a logarithmic transformation of the unique reference reaction count, where we sum the number of coarse-grained and fine-grained references of a candidate reaction:
$p(\texttt{reaction})=\log(n_{\texttt{ref}}(\texttt{reaction})+1)$ (1)
where $n_{\texttt{ref}}(\texttt{reaction})$ represents the unique number of reference reactions in the coarse-grained and fine-grained clusters containing reaction.
3.4 MetaScorer Aggregation
To improve reaction filtering, our MetaScorer integrates scores from Reaction Prior ($s_{\texttt{RP}}$), Reaction Graph Plausibility ($s_{\texttt{RGP}}$), and empirical precedents ($n_{\texttt{ref}}>0$) retrieved via the pipeline in Sec. 3.3. This hybrid approach mitigates the weaknesses of purely data-driven or precedent-based methods. The continuous score is described by equation $s_{\texttt{META}}=\max(s_{\texttt{RGP}},s_{\texttt{RP}})$ if $n_{\texttt{ref}}>0$ (0 otherwise).
For binary classification tasks and search, reactions are filtered using predefined thresholds, which can be selected through grid search to balance precision and recall:
$s_{\texttt{META}}=\begin{cases}1&\text{if }s_{\texttt{RGP}}>\texttt{thr}_{\texttt{RGP}}\text{ and }s_{\texttt{RP}}>\texttt{thr}_{\texttt{RP}}\text{ and }n_{\texttt{ref}}>0\\ 0&\text{otherwise}\end{cases}$ (2)
By synthesizing diverse evidence types MetaScorer enables more reliable reaction filtering for multi-step synthesis planning. This integrated approach mitigates individual weaknesses of purely data-driven or precedent-based methods, yielding improved performance.
3.5 Generator and Integration with Search (Retro*)
For the generator part of RetroTrim, we use the encoder-decoder BART architecture (bart) trained on root-aligned SMILES (zhong2022rootaligned), where the product (target) is processed by the encoder, and the substrates are generated by the decoder. We call this generator RootAligned. The calibrated MetaScorer is used during multi-step retrosynthesis search to improve the quality of the pathways predicted by the BART generator. We integrate the scorer into the Retro* (RetroStar) search framework as a reaction filtering mechanism. Reactions are pruned from the search tree if ${s}_{\texttt{META}}$ defined in 2 is equal to 0.